Two particles are each given a charge of 5 × 10-5 C. What is the force between the charged particles if the distance between them is 2 meters?



1
Expert's answer
2021-05-14T09:41:21-0400

"F=k\frac{q_1q_2}{x^2}=9\cdot10^9\cdot\frac{5\cdot10^{-5}\cdot5\cdot10^{-5}}{2^2}=5.625\ (N)" . Answer